OpenStudy (anonymous):

which of the following is an example of a logarithmic function? y=sin(x) y=4^x y=log10^3 y=x^4

OpenStudy (ikram002p):

its the function that have the variabls on powers or on log . wat u r stuck in ??

OpenStudy (anonymous):

idk how to figure out if it is a logarithmic function

OpenStudy (ikram002p):

ok its logarithmic function when u have y=a^x , such that x is variable . now which one of the options u can ignor ??

OpenStudy (ikram002p):

y=sin(x) is not logarithmic function coz its trigonometric function. y=x^4 is not logarithmic function coz its polynomial

OpenStudy (anonymous):

so would the answer be y=4^x?

OpenStudy (ikram002p):

now u have to chose beween y=4^x y=log10^3 lets take y=log10^3 ,,, ,log10^3=3 y=3 which is constant function (polynomial )

OpenStudy (ikram002p):

u r correct good job :))

OpenStudy (anonymous):

thank you!(:

OpenStudy (ikram002p):

ur wlc.